OpenStudy (anonymous):

What is the ratio of a circle's circumference to it's radius? 1 π/2 √(π) 2π

OpenStudy (compassionate):

π/2, thus it is half the length.

OpenStudy (lgbasallote):

the formula for circumference is \[C = 2\pi r\] does that help?

OpenStudy (anonymous):

and radius =r.

OpenStudy (zzr0ck3r):

the ratio of a circle's circumfrence to it's diameter is pi and radius is (1/2) diamter. hope that helps

OpenStudy (zzr0ck3r):

notice if we have a/b and want (1/2)b then we have a/((1/2)b) = (a/(b/2)) = 2(a/b)

OpenStudy (anonymous):

thank u

OpenStudy (lgbasallote):

@mitchelsewbaran here's another hint... it's not \(\pi /2 \) just so you know lol

OpenStudy (compassionate):

Lol. I just realized my answer is wrong and I still got a medal. YOLO

OpenStudy (anonymous):

ok